Suppose Bull Gator Ben has paid his $14,300 donation and purchased eight season tickets at $32 per game. He discovers that he only needs seven tickets, and sells the eighth ticket for $150.

Considering the donation, how much did Ben pay for each ticket?

Answer is 1819.5

so i am confused how they got it
Answer #1

Total cost of tickets = 14300 + 32 * 8 = 14556

Cost of one ticket paid by Ben = 14556 / 8 = 1819.50